OpenStudy (anonymous):

solve what y is y + 5x = 6;x = -1, 0, 1

OpenStudy (anonymous):

rearrange the equation in terms of y: y= -5x+6 replace x with the given values in the equation: x=-1 y=-5(-1)+6= x=0 y=-5(0)+6= x=1 y=-5(1)+6=

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

that's good:) c ya round!

OpenStudy (anonymous):

c ya